OpenStudy (anonymous):

What is the sum of the roots of the polynomial shown below? F(x)=3x^3+12x^2+3x-18

OpenStudy (butterflydreamer):

sum of roots = -b/a Note that : F(x) = ax^3 + bx^2 + cx + d

OpenStudy (amoodarya):

try to factor (x-1) because x=1 is one of roots (f(1)=0 if you check)

OpenStudy (anonymous):

the sum of the roots are -4

OpenStudy (butterflydreamer):

So you want to find your values of a and b by matching it up :) \[F(x) = ax^3 + bx^2 + cx + d \] \[F(x)= 3x^3 + 12x^2 + 3x - 18\]

OpenStudy (butterflydreamer):

you're close :) just missing a negative sign!

OpenStudy (butterflydreamer):

If you look at the above, we get: a = 3 b = 12 Sum of roots = \[-b/a\] Therefore sum of roots = -12 / 3 = -4

OpenStudy (amoodarya):

he/she did not miss -

OpenStudy (butterflydreamer):

-_- oops, you wrote -4 You're correct! :)

OpenStudy (anonymous):

she lol

OpenStudy (amoodarya):

another observation is to write it \[f(x)=3(x-1)(x+3)(x+2)\]